# RuCCoD: Towards Automated ICD Coding in Russian

### v1

@@ 4.2 Evaluation Methodology
On RuCCoD, our evaluation includes the newly introduced ICD Coding task as well as conventional NER and EL. To recall, ICD Coding can be seen as an entity position-agnostic NER+EL task composition with explicitly removed EHR-level ICD code duplicates. For instance, a language model successfully diagnoses a patient by assigning the correct ICD code when it finds at least one of three mentions of the corresponding ICD disease within an EHR. For all three tasks, we evaluate the standard classification micro-averaged metrics, namely precision, recall, F1-score, and accuracy.
For EL, we use a retrieval-based approach (Liu et al., 2024; Yuan et al., 2022b; Sakhovskiy et al., 2024) and evaluate retrieval accuracy with acc@k, where $acc@k=1$ if a correct ICD code is retrieved at rank $\leq k$. We consider two evaluation scenarios: (i) strict score assessing exact match between a predicted and a ground truth codes and (ii) relaxed score with each code being truncated to its higher-level disease group (e.g., H10.0 mucopurulent is truncated to H10 Conjunctivitis).
@@ 4.3.1 Transfer Learning
First, we performed cross-domain experiments on the entity linking task (with gold entities provided) to see how variability in entities and terminology affect performance. Since UMLS includes the ICD system, we automatically map UMLS CUIs to ICD codes for evaluation. Cross-domain transfer results with entity linking models on RuCCoD, RuCCoN, NEREL-BIO as well as their union are presented in detail in Appx. D. Our evaluation revealed the following key observations.
@@
While extension of ICD vocabulary consistently gives a slightly improved acc@1 in a zero-shot setting, additional synonyms introduce severe noise in a supervised setting. Specifically, a significant drop of 8.1%, 8.4%, and 14.3% acc@1 is observed for SapBERT, CODER, and BERGAMOT, respectively. Even in an unsupervised setting, vocabulary extension drops acc@5 by 5.2% and 6.8% for SapBERT and BERGAMOT, respectively.
Complicated Cross-Terminology Transfer
Both training on RuCCoN and NEREL-BIO as well merge of these corpora with RuCCoD do not lead to improvement over zero-shot coding. The finding indicates that RuCCoN and NEREL-BIO datasets do not easily transfer to our dataset as well as the specificity and high complexity of hierarchical ICD coding within the EL task.
@@
The 15% gap in acc@1 between strict and relaxed evaluations shows the challenging nature of semantically similar diseases within the same therapeutic group.
@@
Regarding the quality of NER across different training sets for disease-related entities, the model trained on the NEREL-BIO dataset and tested on the RuCCoD test set yielded an F1 score of 0.62. The model trained on a combined dataset of NEREL-BIO and RuCCoD achieved scores of 0.72. Similar results were observed with RuCCoN. We also evaluated BINDER, which uses a RuBioBERT backbone and treats NER as a representation learning problem by maximizing similarity between vector representations Zhang et al. . However, BINDER’s performance was 1.5% lower than that of RuBioBERT, which achieved the best F1 score of 0.77 using a softmax classifier. We conclude that the transfer of NER for disease entities is significantly better than for entity linking (EL), with the best results obtained from RuCCoD (full results in Appx. C).
@@ 5.1 Experimental Set-up
Due to highly unbalanced ICD codes distribution, we adopted the F1-score to assess model performance for each class and the weighted F1-score to evaluate overall performance across all classes for the Diagnosis Prediction task. These metrics are well-suited for handling class imbalance, a known challenge in medicine tasks (Johnson and Khoshgoftaar, 2019). For the weighted F1-score, the weight of each class is calculated as the proportion of appointment documents sharing the given ICD code in the union of both training datasets. In our experiments, we evaluate the quality of the models trained on original and linked datasets on the manual test set.
@@ 5.2.1 Diagnosis Prediction Learning
We fine-tuned two Longformer instances: one on the original dataset and one on the linked dataset, for predicting ICD codes from doctor’s appointments. Fig. 5 presents the dependences of the weighted F1-scores on training step number for these two models.
Figure 4: Comparison of weighted F1 scores on the manual diagnosis prediction test set for models trained on original and linked datasets at different training steps.
@@
As seen from Fig. 5, AI-guided ICD coding (linked data) significantly outperforms manual coding (original data) with the peak weighted F1-score of 0.48. The latter quickly reaches its F1-score plateau at 0.2. The huge performance gap of 0.28 highlights the effectiveness of automatic data annotation for model training. Yet, the finding reveals the complexity of ICD-agreed diagnosis prediction task for professional physicians indicating the necessity of AI-driven assistance.
@@ 5.3 Diagnosing Stability to Disease Frequency
The test dataset was split into two parts: the 10% most frequent ICD codes and the 10% least frequent ICD codes, with a minimum frequency threshold of 15 instances in the manual test set for the less frequent group. The stratification approach is designed to align with the distribution of real-world diagnoses assigned and carefully verified by clinicians.
@@ 5.4 Disease-Wise Quality Shift Analysis
Fig. 6 displays how changes in appointment counts from original to linked data affect the diagnosis prediction F1 score. Notably, F1 scores generally improved across the majority of ICD codes regardless of whether automatic linking increased or decreased the number of appointments. This suggests improved class balance in the linked dataset, although the effect varies.
@@
In Fig. 6, a sharp F1 score drop is observed for I25.2 Past myocardial infarction. Apparently, the disease has been mistakingly re-linked to other errouness by our ICD Coding system. We studied the case by analyzing which ICD codes has I25.2 been replaced with. As shown in Fig. 7, the most frequent transition, from I25.2 to I11.9 (Hypertensive heart disease, 29,639 cases), resulted in a minimal F1 improvement of 0.02, likely due to overlapping symptoms. The transition to E11.9 (Type 2 diabetes mellitus, 11,413 cases) produced the highest F1 gain of 0.48, reflecting clearer distinctions. Transitions to I25.1 (Atherosclerotic heart disease, 12,932 cases) and I20.9 (Angina pectoris, 11,819 cases) led to significant F1 increases of 0.38 and 0.47, while I67.9 (Unspecified cerebrovascular disease, 10,573 cases) showed a moderate gain of 0.21. These results suggest that diagnoses with clearer distinctions improve F1 more than those with overlapping symptoms.
@@ 6 Conclusion
In this paper, we presented the first models for multi-label ICD-10 coding of electronic health records (EHRs) in Russian. Our study focuses on two key tasks: information extraction from the diagnosis field of EHRs and diagnosis prediction based on a patient’s medical history. The NLP pipeline developed for the first task was utilized to re-annotate EHRs in the training set for the second task. The results demonstrate that fine-tuned LMs significantly enhance performance in predicting ICD codes from past medical history. Specifically, the model trained on automatically linked data exhibited faster learning and better generalization compared to the original dataset, achieving higher weighted F1 scores early in training, while the original model plateaued with minimal improvements. Notably, the linked data model consistently outperformed the original across both frequent and rare ICD classes, achieving higher F1 scores with reduced variability. This suggests that the linked dataset enables effective handling of both common and rare ICD codes. Overall, our findings highlight the importance of a neural pipeline for automating ICD coding and improving the accuracy and informativeness of medical text labeling.
@@
The most relevant corpora to our study are RuCCoN Nesterov et al. (2022b) and NEREL-BIO Loukachevitch et al. (2023, 2024), which link entities from clinical records or PubMed abstracts to the Russian segment of UMLS. We conducted preliminary transfer learning experiments using these two corpora; however, a detailed analysis of the semantic differences among the three corpora has yet to be performed.
Moderate Inter-Annotator Agreement
Among ICD codes, the IAA value was 50%, indicating moderate agreement, while for ICD groups, the IAA increased to 74%, reflecting higher consistency at a group-wise level. This disparity suggests that annotators were generally aligned when categorizing broader ICD groups but faced challenges in granular code assignment. This pattern mirrors trends observed in clinical practice, possibly due to ambiguities in documentation and coding guidelines (cf. §3.1). While our terminal code IAA (50%) aligns with the upper bounds of reported expert agreement (29.2%–46.8%) Stausberg et al. (2008), the residual variability underscores the need for standardized annotation protocols or ensemble approaches to mitigate subjectivity in fine-grained coding.
@@
While our dataset is substantial, it may not fully capture the diversity of clinical scenarios and patient demographics. A more varied dataset could improve the robustness and generalizability of the models. Clinical language can vary significantly across different medical specialties and institutions. This variability may impact the model’s ability to generalize across various clinical contexts.
@@
The dataset may suffer from class imbalance, with certain ICD codes being underrepresented. This could affect the model’s ability to generalize and accurately predict less common diagnoses.
@@
RuCCoD does not contain any personally identifiable patient information. The dataset consists solely of diagnosis conclusions written by medical professionals, which were manually labeled based on the ICD-10 CM (Clinical Modification) system. Prior to the annotation process, annotators were instructed to ensure that no personal information was included in the conclusions. Their task was to identify and remove any personal or identifiable information manually from these texts. Overall, no patient-related information will be disclosed upon the dataset’s release.
@@
The dataset introduced in this paper involved only new annotations. Dataset annotation was conducted by annotators, and there are no associated concerns (e.g. regarding compensation). Each annotator received a compensation of approximately \$12 per hour for their contributions. An estimated 85 hours of annotation work per expert resulted in a total payment of \$1,020 per annotator. For context, the minimum monthly wage in Russia for full-time employment is under \$200, highlighting the substantial effort and investment in creating this high-quality resource. All annotators were aware of potential annotation usage for research purposes. As discussed in limitations, we believe these new annotated datasets serve as a starting point for the evaluation of LMs on ICD coding in Russian. Our annotations, code, and annotation guidelines will be released upon acceptance of this paper.
@@
The RuCCoD and all associated annotation materials are being released under the CC BY 4.0 license. It should be noted that the dataset contains only diagnosis codes and no medical histories or personal patient data. Furthermore, all diagnoses have been rigorously verified to ensure complete anonymity, in accordance with the prevailing norms of open research practice. Our GitHub repository and HuggingFace dataset card will include comprehensive documentation on the codebase, the methodology for creating benchmarks, and the human annotation process. The source code for our experiments will be freely available at this anonymized repository: https://github.com/auto-icd-coding/ruccod.
@@
We utilize Grammarly to enhance and proofread the text of this paper, correcting grammatical, spelling, and stylistic errors, as well as rephrasing sentences. Consequently, certain sections of our publication may be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
Appendix A Appendix: Annotation guidelines.
@@ Appendix B Related Work
Klang et al. Klang et al. (2024) demonstrated the effectiveness of RAG in enhancing LLMs for ICD-10-CM medical coding. Their study revealed that RAG-enhanced LLMs outperform human coders in accuracy and specificity, emphasizing the potential of retrieval mechanisms in improving clinical documentation. Similarly, Kwan Kwan (2024) proposed a two-stage Retrieve-Rank system for medical coding, achieving a perfect match rate for ICD-10-CM codes and significantly surpassing vanilla LLMs. The MedCodER framework Baksi et al. (2024) leverages a pipeline of extraction, retrieval, and reranking, to improve automation and interpretability in ICD-10 coding. It demonstrates SOTA performance on ACI-BENCH by integrating LLMs with semantic search and evidence-based reasoning. Boyle et al. Boyle et al. (2023) presented a zero-shot ICD coding approach using LLMs and a tree-search strategy, achieving a SOTA on the CodiEsp dataset, particularly excelling in rare code prediction without task-specific training. Abdulnazar et al. Abdulnazar et al. (2024) applied GPT-4 for clinical text cleansing to enhance MCN. By combining text standardization with RAG, their method improved mapping precision to SNOMED CT in the German language.

### v2

@@ 4.2 Evaluation Methodology
On RuCCoD, our evaluation includes conventional NER and EL as well as end-to-end document-level code assignment with EHR-level code aggregation (Sec. 2). To recall, document-level metrics is an entity position-agnostic NER+EL task composition with explicitly removed EHR-level ICD code duplicates. For instance, a language model successfully diagnoses a patient by assigning the correct ICD code when it finds at least one of three mentions of the corresponding ICD disease within an EHR. For all three tasks, we report accuracy and the micro-averaged precision, recall, F1-score.
Following prior research (Phan et al., 2019; Wright et al., 2019; Liu et al., 2021a; Yuan et al., 2022b; Sakhovskiy et al., 2024), we use a retrieval-based EL approach and evaluate retrieval accuracy: $acc@k=1$ if a correct ICD code is retrieved at rank $\leq k$. We consider two evaluation scenarios: (i) strict score assessing exact match between a predicted ground truth codes; (ii) relaxed score with each code being truncated to higher-level disease group (e.g., H10.0 Mucopurulent conjunctivitis is truncated to H10 Conjunctivitis).
@@ 4.3.1 Transfer Learning
First, we performed cross-domain experiments on EL to see how variability in entities and terminology affects the performance. Since UMLS includes the ICD system, we automatically map UMLS CUIs to ICD codes for evaluation. Cross-domain transfer results with entity linking models on RuCCoD, RuCCoN, NEREL-BIO and their union are presented in detail in Appx. D. The evaluation has revealed the following key observations.
@@
While extension of ICD vocabulary consistently gives a slightly improved acc@1 in a zero-shot setting, additional synonyms introduce severe noise in a supervised setting. Specifically, a significant drop of 8.1%, 8.4%, and 14.3% acc@1 is observed for SapBERT, CODER, and BERGAMOT, respectively. Even in an unsupervised setting, vocabulary extension drops acc@5 by 5.2% and 6.8% for SapBERT and BERGAMOT, respectively. The results highlight the specifity of ICD terminology within the medical domain.
Limitations of Cross-Terminology Transfer
Both training on RuCCoN and NEREL-BIO as well as merge of these corpora with RuCCoD do not lead to improvement over zero-shot coding (Tab. 7, Appx. D). Thus, training on other EL corpora normalized to the UMLS vocabulary does not transfer to our RuCCoD dataset normalized to ICD vocabulary. The finding indicates the specificity and high complexity of hierarchical ICD coding within the EL task.
@@
The 15% gap in acc@1 between the strict and relaxed evaluation (Tab. 7, Appx. D) shows the challenging nature of semantically similar diseases within the same therapeutic group. The finding further illustrates the underlying challenge for human ICD coders: correct disease specification with an appropriate code is a far more complex task than general disease identification.
@@
A NER model trained on the disease-related entities from NEREL-BIO gained an F1 score of 0.62 on RuCCoD’s test set. The model trained on a combined dataset of NEREL-BIO and RuCCoD achieved scores of 0.72. Similar results were observed with RuCCoN. We also evaluated BINDER, which uses a RuBioBERT backbone and treats NER as a representation learning problem by maximizing similarity between vector representations Zhang et al. . However, BINDER’s performance was 1.5% lower than RuBioBERT’s, which gained the best F1 score of 0.77 with a softmax classifier. NER transfer for disease entities is significantly better than for entity linking (EL), with the best results obtained from RuCCoD (full results are in Appx. C).
@@ 4.3.2 End-to-end ICD coding
Figure 4: Comparison of weighted F1 scores on the common test set for models trained on original and linked datasets at different training steps.
@@ 5 Diagnosis Prediction Evaluation
For the Diagnosis Prediction task, we explore whether automatic relabeling of ICD codes assigned to an EHR increases the robustness and reliance of a diagnostic model. To address this, we train two diagnostic models on RuCCoD-DP’s (i) original and (ii) linked independently. These two models learn from the same EHRs but use different target ICD labels (doctor-assigned and AI-based codes, respectively). Given that both annotation types are produced from human-written diagnosis conclusions, our experimental design compares two methodologies for formalizing a verbose diagnostic conclusion into a set of ICD codes.
@@ 5.1 Experimental Set-up
In our experiments, we evaluate the quality of the models trained on the original and linked datasets on the common test set. Because training datasets contain a much larger and more diverse set of ICD codes than the test set, all evaluation metrics are computed only over the codes in the test set to ensure a consistent and comparable label space.
To address the imbalanced long-tail ICD code distribution in the Diagnosis Prediction task, we adopt the weighted F1 score for overall evaluation, a standard choice in prior work on automated ICD coding (Johnson and Khoshgoftaar, 2019; Blanco et al., 2020). The weight of each class is calculated as the proportion of EHRs sharing the given ICD code in the union of both training datasets. Per-class F1 scores were also measured to explore performance variations across frequent and rare ICD codes.
Beyond the weighted F1, we additionally report micro-averaged confusion components—True Positives (TP), False Positives (FP), False Negatives (FN), and True Negatives (TN). In the multi-label setting, each (record, code) pair is treated as a binary outcome, and these components provide a more granular view of diagnostic sensitivity and specificity than F1 alone.
@@ 5.2.1 Diagnosis Prediction Learning
To predict ICD codes from doctor’s appointments, we fine-tuned two Longformer models, one using the original dataset and the other using the linked dataset. The weighted F1-scores for the two models against the training count are shown in Fig. 4.
@@
As seen from Fig. 4, the model trained on AI codes coding (linked data) significantly outperforms the one trained on manual codes (original data) with the peak weighted F1-score of 0.48. The latter quickly reaches its F1-score plateau at 0.2. The huge performance gap of 0.28 highlights the effectiveness of automatic data annotation for model training. The finding reveals the complexity of the ICD-agreed diagnosis prediction task for professional physicians, indicating the need for AI-driven assistance. For instance, an AI diagnosis conclusion can serve as an independent second opinion, while the final diagnosis should be determined by a qualified human expert.
@@ 5.3 Diagnosing Stability to Disease Frequency
The test dataset was split into two parts: the 10% most frequent ICD codes and the 10% least frequent ICD codes, with a minimum frequency threshold of 15 instances in the common test set for the less frequent group. The stratification approach is designed to align with the distribution of real-world diagnoses assigned and carefully verified by clinicians.
@@
Top 10% codes Bottom 10% codes
Dataset FN FP TN TP FN FP TN TP
Original 194 146 967 175 36 630 1284 26
Linked 118 136 977 251 17 57 1857 45
Table 4: Micro-averaged confusion components on the common test set for the 10% most frequent (Top) and 10% least frequent (Bottom) ICD codes, depending on whether the model is trained on original or linked data.
We further decompose the results into confusion components to clarify how code re-linking affects different frequency groups. As seen from Table 4, training on linked data reduces False Negatives by 40% and improves True Positives by 43% for the most frequent codes compared with the training on the original data. For the least frequent codes, we observe a 53% decrease in False Negatives and a 91% decrease in False Positives, while True Negatives and True Positives increase by 45% and 73%, respectively. This analysis confirms that the improvements from linked training align with the F1 gains, reflecting fewer missed and spurious predictions and more correct detections across both frequent and rare codes.
@@ 5.4 Disease-Wise Quality Shift Analysis
Fig. 6 shows how changes in appointment counts from original to linked data affect the diagnosing F1 score. Notably, F1 scores generally improved for the majority of ICD codes regardless of appointment counts increase or decrease. Although the performance change varies, the observations suggest the improved class balance. The only exception is the I25.2 Past myocardial infarction code, which will be analyzed in detail in the subsequent section.
@@
In Fig. 6, a sharp F1 score drop is observed for I25.2 Past myocardial infarction. Apparently, the disease has been mistakingly re-linked to other errounessly. We studied the case by analyzing which ICD codes has I25.2 been replaced with. Fig. 7 shows the most frequent transition (I25.2 to I11.9, Hypertensive heart disease) yielded minimal F1 improvement (0.02), likely due to symptom overlap. E11.9 (Type 2 diabetes mellitus) had the highest gain (0.48) due to clearer distinctions. I25.1 (Atherosclerotic heart disease) and I20.9 (Angina pectoris) had significant gains (0.38, 0.47), while I67.9 (Unspecified cerebrovascular disease) had a moderate gain (0.21). From the results, clearly distinguishable diagnoses yield higher F1 scores compared to those with symptom overlap.
@@ 6 Conclusion
In this paper, we presented the first models for multi-label ICD-10 coding of electronic health records (EHRs) in Russian. Our study focuses on two key tasks: information extraction from the diagnosis field of EHRs and diagnosis prediction based on a patient’s medical history. The NLP pipeline developed for the first task was utilized to re-annotate EHRs in the training set for the second task. The results demonstrate that fine-tuned LMs significantly enhance performance in predicting ICD codes from past medical history. Specifically, the model trained on automatically linked data exhibited faster learning and better generalization compared to the original dataset, achieving higher weighted F1 scores early in training, while the original model plateaued with minimal improvements. Notably, the linked data model consistently outperformed the original across both frequent and rare ICD classes, achieving higher F1 scores together with improved confusion metrics (TP/FP/FN/TN) and reduced variability. These results suggest that the linked dataset enables effective handling of both common and rare ICD codes. Overall, our findings highlight the importance of a neural pipeline for automating ICD coding and improving the accuracy and informativeness of medical text labeling.
@@
The most relevant corpora to our study are RuCCoN Nesterov et al. (2022) and NEREL-BIO Loukachevitch et al. (2023, 2024), which link entities from clinical records or PubMed abstracts to the Russian part of UMLS. We performed preliminary transfer learning experiments with these two corpora, but a thorough analysis of the semantic differences across all three corpora is future work.
@@
Our dataset may not fully capture the diversity of clinical scenarios and patient demographics. A more varied dataset could improve the robustness and generalizability of the models. Clinical language can vary significantly across different medical specialties and institutions. This variability may impact the model’s ability to generalize across various clinical contexts.
@@
The dataset may suffer from class imbalance, with certain ICD codes being underrepresented. This could affect a model’s ability to accurately predict less common diagnoses.
@@
RuCCoD does not contain any personally identifiable patient information. The dataset consists solely of diagnosis conclusions written by medical professionals, which were manually labeled based on the ICD-10 CM (Clinical Modification) system. Prior to the annotation process, annotators were instructed to ensure that no personal information was included in the conclusions. Their task was to identify and remove any personal or identifiable information manually from these texts. No patient-related information is disclosed in RuCCoD.
@@
All data annotations in this paper are newly created. Dataset annotation was conducted by annotators, and there are no associated concerns (e.g. regarding compensation). Each annotator was paid a rate of \$12 per hour for their contributions. An estimated 85 hours of annotation work per expert resulted in a total payment of \$1,020 per annotator. Russia’s full-time monthly minimum wage is under \$200, highlighting the substantial effort and investment in creating this high-quality resource. All annotators were aware of potential annotation usage for research purposes.
@@
The RuCCoD and all associated annotation materials are being released under the CC BY 4.0 license. It should be noted that the dataset contains only diagnosis codes and no medical histories or personal patient data. Furthermore, all diagnoses have been rigorously verified to ensure complete anonymity, in accordance with the prevailing norms of open research practice.
@@
We used Grammarly to proofread the paper, correcting grammatical, spelling, and stylistic errors, as well as rephrasing sentences. Consequently, certain sections of our paper may be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
Acknowledgements
This work was supported by a grant, provided by the Ministry of Economic Development of the Russian Federation in accordance with the subsidy agreement (agreement identifier  [^00] C313925P4G0002) and the agreement with the Ivannikov Institute for System Programming of the Russian Academy of Sciences dated June 20, 2025 No. 139-15-2025-011.
Appendix A Appendix: Annotation Details
@@ A.4 Annotation Tool
A.5 Inter-Annotator Agreement
To assess consistency among experts manual ICD coding, we first measured Inter-Annotator Agreement (IAA) metric, defined as the ratio of accepted codes to the total number of unique codes assigned per record Luo et al. (2019). The overall IAA reached 50%, indicating a moderate level of agreement. Such values align with previous reports on ICD coding, where Kappa coefficients typically range from 27% to 42% (corresponding to agreement rates of 29.2%–46.8%) Stausberg et al. (2008), and coding accuracy varies widely from 41.8% to 98% depending on domain and code granularity Campbell et al. (2001); Hosseini et al. (2021). These observations confirm that moderate agreement is expected given the inherent complexity and subjectivity of assigning fine-grained ICD codes.
To further ensure the consistency of obtained annotations and provide an additional comparative assessment of agreement, we also measured agreement using the Jaccard similarity coefficient (JSC). This metric captures the ratio of the intersection to the union of ICD code sets assigned by two annotators and is particularly suited for multi-label settings where overlapping but not identical code sets are common. The mean pairwise JSC values were 0.308, 0.336, and 0.366 for annotator pairs (1–2), (1–3), and (2–3), respectively. These results are consistent with existing studies on ICD coding from unstructured clinical text, which report average Jaccard agreements ranging from 0.179 to 0.345 and, for certain code categories, scores approaching zero Xu et al. (2019); Lucas et al. (2021).
Together, the IAA and JSC analyses indicate that the observed annotation variability is realistic and reflects the intrinsic difficulty of achieving high inter-expert consistency in multi-label ICD coding.
@@ Appendix B Related Work
In describing our work, we encountered persistent terminological ambiguity arising from overlapping nomenclature for distinct task formulations. For instance, the term “ICD coding” is broadly applied to both (1) multi-label classification of medical texts (e.g., assigning ICD codes to discharge summaries) Li and Yu (2020); Vu et al. (2020); Wang et al. (2024) and (2) entity linking, where discrete clinical diagnoses are mapped to specific codes Lavergne et al. (2016); Névéol et al. (2017); Coutinho and Martins (2022). This conflation obscures fundamental differences: the former treats coding as document-level prediction to capture all relevant codes for a patient’s condition, while the latter focuses on precise alignment of clinical entities (e.g., distinguishing “acute myocardial infarction” from its subtypes) through semantic matching, addressing challenges like synonymy or hierarchical code relationships. To resolve this ambiguity, in our work we propose explicit terminology: “ICD coding” refers to multi-label classification of medical texts, whereas “Medical entity linking” denotes entity-level code assignment.
@@
Klang et al. (2024) demonstrated the effectiveness of RAG in enhancing LLMs for ICD-10-CM medical coding. Their study revealed that RAG-enhanced LLMs outperform human coders in accuracy and specificity, emphasizing the potential of retrieval mechanisms in improving clinical documentation. Similarly, Kwan (2024) proposed a two-stage Retrieve-Rank system for medical coding, achieving a perfect match rate for ICD-10-CM codes and significantly surpassing vanilla LLMs. The MedCodER framework Baksi et al. (2024) leverages a pipeline of extraction, retrieval, and reranking, to improve automation and interpretability in ICD-10 coding. It demonstrates SOTA performance on ACI-BENCH by integrating LLMs with semantic search and evidence-based reasoning. Boyle et al. Boyle et al. (2023) presented a zero-shot ICD coding approach using LLMs and a tree-search strategy, achieving a SOTA on the CodiEsp dataset, particularly excelling in rare code prediction without task-specific training. Abdulnazar et al. Abdulnazar et al. (2024) applied GPT-4 for clinical text cleansing to enhance MCN. By combining text standardization with RAG, their method improved mapping precision to SNOMED CT in the German language.